# line of sight CSI channel state information

### v2

@@ VI Key FR3 features in $6$G Standards
  • Multi-band pilot design, which aims at designing hybrid pilot structures that combine common and band-specific pilots while employing sparse allocation, frequency sharing, and adaptive density. Multi-Band pilot design can be realized via AI-driven optimization and compressed sensing techniques to minimize redundancy and intelligently manage pilots for greater efficiency in multi-band environments.
  • Agile timing advance, which aims at developing new timing advance strategies to ensure seamless switching and synchronization across carriers in different frequency bands, which is critical for achieving low-latency and high-reliability performance in $6$G.
  • Multi-carrier synchronization and scalable numerology, meaning that it is vital to introduce enhanced scalable numerology to support multi-carrier synchronization across diverse FR3 bandwidths, which further allows adaptive subcarrier spacing and multi-band coordination without excessive signaling overhead.
[…]
From RAN perspective, the open radio access network (O-RAN) RAN intelligent controller (RIC) plays a crucial role in enabling spectral agility and intelligence by leveraging real-time (RT) and non-RT network optimizations. Through the use of xApps (near-RT RIC applications) and rApps (Non-RT RIC applications), the RIC can dynamically manage spectrum resources, adapt to changing conditions, and optimize performance in FR3. Hereafter, we examine how spectral agility and spectral intelligence can be achieved through the RIC.
Spectral Agility: The ability of a network to dynamically reconfigure spectrum usage based on real-time conditions, interference levels, traffic demands, and regulatory constraints can be achieved through the following mechanisms:
  • Dynamic carrier aggregation & multi-band adaptation: The Near-RT RIC can monitor real-time spectrum conditions and trigger carrier aggregation adjustments across multiple FR3 bands, which enables on-the-fly switching between licensed, shared, and unlicensed spectrum resources, optimizing throughput and reducing congestion.
  • Intelligent spectrum sensing and sharing: By integrating AI/ML models within rApps, the Non-RT RIC can analyze historical spectrum usage patterns and proactively allocate spectrum to different cells or operators, thus enabling dynamic spectrum access and coexistence with satellite, radar, or other wireless services operating in FR3.
Spectral Intelligence: Efficiency, reliability, and adaptability of spectrum can be achieved through intelligent decision-making and automation.
  • Beamforming, channel estimation, and interference management: xApps in the Near-RT RIC continuously monitor spectrum conditions and adjust beamforming strategies for multi-user MIMO in FR3. Advanced channel estimation and interference prediction can help minimize co-channel and adjacent-band interference, ensuring optimal signal quality.
  • Multi-band O-RU control for distributed MIMO: The Near-RT RIC can coordinate multi-band O-RUs across FR3 to enable distributed MIMO architectures. Leveraging O-RAN’s open fronthaul interface, it ensures tight synchronization across spatially separated RUs, maximizing spectral efficiency.
  • Power efficiency: Other key innovations that can be implemented by leveraging the RIC are adaptive sampling and dynamic range control. By adjusting these parameters based on real-time spectral conditions, FR3 O-RUs reduce unnecessary power consumption and enhance operational efficiency.
VII Conclusions
@@
This paper outlined a vision for $6$G in the FR3 spectrum, exploring its upper mid-band potential, key challenges, and necessary standardization changes. We highlighted FR3’s advantages in capacity, coverage, and SE, while emphasizing the impact of its channel characteristics on beam management. Based on our FR3 measurements, we underscored the need for frequency hopping to mitigate blockages, maximize SE, and incumbent coexistence, reinforcing the importance of spectrum agility. We examined the potential of mMIMO for FR3 and discussed design challenges, particularly in the uplink. For sensing, we analyzed performance bounds for FR3 ISAC multiband sensing. Additionally, we detailed key $6$G standardization features, including 3GPP radio frame modifications and O-RAN-based innovations, using RIC-driven xApps and rApps for intelligent spectrum management. However, open questions remain regarding FR3 propagation characteristics, PHY design, and regulatory considerations.

### v3

@@ V Open Questions & Potential Solutions
  • Multi-band pilot design, which compresses and shares pilots intelligently via sparse allocation, in addition to frequency sharing and adaptive density. Multi-Band pilot design can be realized via AI-driven optimization and compressed sensing techniques to minimize redundancy and manage pilots for better efficiency in multi-band environments.
  • Agile timing advance and scalable numerology, which keep devices tightly synchronized as they hop among carriers with different sub-carrier spacings, sustaining low-latency links.
In addition, an AI-guided RAN intelligent controller (RIC) orchestrates carrier aggregation, beamforming, interference mitigation, and power-adaptive O-RUs so the full FR3 spectrum is exploited efficiently and energy-consciously. Near-RT xApps run inside the RIC’s real-time loop, reacting within milliseconds to spectrum-usage shifts to retune carriers and beams, giving the network reflex-like agility.
Regarding legacy RAN, even where incumbent non-O-RAN platforms dominate, 66G spectrum agility can still be achieved. Operators can for example: $(i)$ retrofit wideband and multiband remote radio units and software-defined basebands already supported by most LTE and 55G sites; $(ii)$ activate standards-based features such as DSS, expanded carrier aggregation, in addition to integration with other important technologies such as self-organizing networks and auto refarming. $(iii)$ In addition, an AI-driven, closed-loop assurance layer that ingests real time RAN data, processes it through a machine-learning pipeline, and then issues automated network actions allows carriers to be split, resized, or retuned on the fly whenever slice traffic or service requirements change. Such approaches can deliver O-RAN like features without a disruptive replacement of legacy RAN hardware.
VI Conclusions
@@
This paper outlined a vision for 66G in the FR3 spectrum, exploring its upper mid-band potential, key challenges, and necessary standardization changes. We highlighted FR3’s advantages in capacity, coverage, and SE, while emphasizing the impact of its channel characteristics on beam management. Based on our FR3 measurements, we underscored the need for frequency hopping to mitigate blockages, maximize SE, and incumbent coexistence, reinforcing the importance of spectrum agility. We examined the potential of mMIMO for FR3 and discussed design challenges, particularly in UL. For sensing, we analyzed performance bounds for FR3 ISAC multiband sensing. Additionally, we detailed key 66G standardization features, including 3GPP radio frame modifications and O-RAN-based innovations, using RIC-driven xApps and rApps for intelligent spectrum management. However, open questions remain regarding FR3 propagation characteristics, PHY design, and regulatory considerations.